crispyrat:

Let ` r ` and ` s ` be the roots of \(3x^2 + 4x + 12 = 0\) Find \(r^2 + s^2\)

crispyrat:

i know i have to do the viesta laws so sum of r and s is -4/3 and the product is 4

snowflake0531:

What are those dollar signs-

crispyrat:

it was for latex but i didn't work

snowflake0531:

Can you retype the question

crispyrat:

k Let r and s be the roots of 3x^2 + 4x + 12 = 0. Find r^2 + s^2.

snowflake0531:

Can you find the roots of \[3x^2 +4x+12\] For example, since you can't do grouping or factoring in this, can you use the quadratic formula to find the roots?

crispyrat:

yee im having trouble with that so idk :( like i can find the sum and product but thats about it

snowflake0531:

Do you know what the quadratic formula is

crispyrat:

yes its ax^2+ax+bx+c

snowflake0531:

No- \[ax^2 + bx+c\] is standard form for a quadratic equation But the quadratic formula is \[\frac{ (-b \pm \sqrt(b^2 -4ac)) }{ 2a }\]

crispyrat:

oh cool what can i use that for?

snowflake0531:

ye so, we can plug in the a, b, and c a=3 b=4 c=12 Can you plug in these values into that

snowflake0531:

i need to go-

snowflake0531:

kek so you would do \[\frac{ -4\pm \sqrt(16-4(3)(12)) }{ 6 }\] which condenses to \[\frac{ -4\pm \sqrt(-128) }{ 6 }\] can you simplify that?._.

crispyrat:

im sorry i don't understand. i thout square roots couldnt be negative?

snowflake0531:

They can't, that's why they turn into i i is \[\sqrt-1\]

crispyrat:

im sorry we didn't learn this in my class ;-;

snowflake0531:

well, just disregard the negative for now, how would you simplify it

Angle:

you don't need to worry about the negative in the square root because the r^2 and s^2 you are looking for are going to get rid of the square roots later anyways

crispyrat:

so -4*sqrt(128 is the same as -4*8sqrt(2) which is -32sqrt(2) ?

Angle:

wait, why multiply the -4?

Angle:

example: \(2 \pm3 = 5 ~or~-1\)

crispyrat:

of so its adding my bad

snowflake0531:

It's adding adn subtracting And then it's a _ or _ as answer

snowflake0531:

Just disregard that symbol for now too, xd Let's just first simplify the sqrt(128)

crispyrat:

its 8sqrt(2)

snowflake0531:

Yea So, we then get \[\frac{ -4\pm8\sqrt-2 }{ 6 }\] Can you now simplify the, 4, 8, 6 stuff

snowflake0531:

Just divide the numerator and denominator by 2

crispyrat:

-2+/-4sqrt(-2)/3

Angle:

yus this means that \(\large\frac{-2+4\sqrt{-2}}{3}\) is one "root" and the other root is \(\large\frac{-2-4\sqrt{-2}}{3}\)

crispyrat:

oh how would i apply r^2+s^2 to that tho

crispyrat:

like i guess i plug it in but im a bit confused on how to calculate

Angle:

\(\huge(\frac{-2+4\sqrt{-2}}{3})^2+(\frac{-2-4\sqrt{-2}}{3})^2\)

crispyrat:

i'm sorry i don't understand on how to calculate that :(

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle \(\huge(\frac{-2+4\sqrt{-2}}{3})^2+(\frac{-2-4\sqrt{-2}}{3})^2\) \(\color{#0cbb34}{\text{End of Quote}}\) \(\large(\frac{-2+4\sqrt{-2}}{3} \times \frac{-2+4\sqrt{-2}}{3})+(\frac{-2-4\sqrt{-2}}{3}\times \frac{-2-4\sqrt{-2}}{3})\)

crispyrat:

ok so would it be (-2+8)/3^2=6/3^2=2^2=4 and also (-2-8/3)^2=-10/3=-100/9 did i do something wrong???

Angle:

\((-2+4\sqrt{-2})(-2+4\sqrt{-2}) = 4 + 8\sqrt{-2}+8\sqrt{-2}+16(-2)\)

crispyrat:

-32+4+16sqrt(-2)=-28+16sqrt(-2)

Angle:

yes and do the same with the other set

crispyrat:

i got 4-8sqrt(-2)-8sqrt(-2)+8=4-8-16sqrt(2)=-4-16sqrt(-2)

crispyrat:

i may have made a typo

crispyrat:

is it 4+8sqrt(-2)+8sqrt(-2)+8 and 12+8sqrt(-2)

Angle:

4-8sqrt(-2)-8sqrt(-2)+16(-2) = 4 - 16sqrt(-2) - 32

crispyrat:

then i add those togheter 4 - 16sqrt(-2) - 32+-28+16sqrt(-2)=-56

crispyrat:

and so uts -56/3? or -18 2/3

Angle:

over 9?

crispyrat:

yep i got it write thx so much ^-^

Angle:

:)